# A review on combinatorial approach to aggregation

### v1

@@ II.8 Generating function method
Below, we will show how to obtain Eq. (11) from Eq. (12). This derivation Lepek_2019 is worth reviewing because of (at least) two important reasons. First, it proves that Eqs. (11) and (12) define the same sequence. The recursive form will be later used to obtain $x_{g}$ for any given kernel. Second, the methodology to obtain Eq. (11) from Eq. (12), i.e. generating function method, was the main mathematical tool applied to derive explicit solutions to several important kernels (e.g., product, additive, condensation, electrorheological).
Expanding binomials we have
@@
Transforming left–hand side gives
@@
$G\left(z\right)\equiv\sum^{\infty}_{g=1}{y_{g}z^{g}}$ (21)
is the generating function for $y_{g}$. Analogously, after transforming right–hand side of Eq. (19) we obtain
@@
By the above transformations we obtained differential equation for $G\left(z\right)$,
$z\frac{\partial}{\partial z}G\left(z\right)-G\left(z\right)=\frac{1}{2}{\left(G\left(z\right)\right)}^{2}.$ (23)
As Eq. (23) is an ODE with separated variables, its solution may be written as
$\begin{split}G\left(z\right)&=\frac{2Az}{1-Az}=2Az\left(\frac{1}{1-Az}\right)\\ &=2Az\sum^{\infty}_{g=0}{{\left(Az\right)}^{g}}=2\sum^{\infty}_{g=1}{{\left(Az\right)}^{g}}\end{split}$ (24)
with $A$ being a constant.
@@
$2A^{g}=\frac{x_{g}}{g!\left(g-1\right)!}.$ (25)
Assuming reasonably that $x_{1}=1$, we can calculate the constant $A=1/2$ to obtain
@@
II.9 Expressions for average number of clusters of given size and its standard deviation
Armed with the methodology to calculate $x_{g}$, we can proceed to find an average number of clusters of a given size, taking advantage of Eq. (7) Fronczak_2018 ; Lepek_2019 .
First, we will write down the expressions for the thermodynamic probabilities $W(\mathrm{\Omega})$,
$\begin{split}W(\mathrm{\Omega})&=\left[t!\prod^{N-k+1}_{g=1}{\frac{1}{{\left(\left(g-1\right)!\right)}^{n_{g}}}}\right]\left[N!\prod^{N-k+1}_{g=1}{\frac{1}{n_{g}!}{\left(\frac{x_{g}}{g!}\right)}^{n_{g}}}\right]\\ &=t!N!\prod^{N-k+1}_{g=1}{\frac{1}{n_{g}!}{\left(\frac{x_{g}}{(g-1)!g!}\right)}^{n_{g}}},\end{split}$ (27)
and for the normalizing factor $Z$ calculated over all states of the system,
where, for further simplification, we used ${\omega}_{g}=x_{g}/(g-1)!$.
Considering the above, the probability distribution is specified as
Eq. (30) was the main contribution of Ref. Fronczak_2018 as it provides the most detailed information about the finite–size coalescing system. Let us recall that t=Nkt=N-k; thus, Eq. (30) holds for any moment of the evolution.
where, in the last transformation, we use expression for the derivative of the Bell polynomial (described in Appendix A, Eq. (116)) and we use the following substitutions consistently,
$\omega_{s}=\frac{x_{s}}{(s-1)!}\;\;\;\;\textrm{and}\;\;\;\;\left\{\omega_{g}\right\}=\left\{\frac{x_{g}}{(g-1)!}\right\}.$ (36)
Finally, we obtain (cf. Figs. 6 and 7)
but we will use the form with $k$ for shorter notation.
It should also be noted that $\omega_{s}$ is a single value depending on the cluster size $s$ while $\omega_{g}$ is a sequence which is independent of $s$, where $g$ varies from $1$ to Nk+1N-k+1 (i.e., to $t+1$), see Sec. II.4. Clearly, for cluster sizes $s\geq t+2$ we have $\left\langle n_{s}\right\rangle=0$ as clusters of such sizes had no chance to arise.
A significant element that makes the combinatorial approach different from the previous approaches to coagulation (Smoluchowski, Marcus–Lushnikov) is the fact that using Eq. (30), one can calculate standard deviation corresponding to the average number of clusters of a given size as Fronczak_2018
@@ II.9 Expressions for average number of clusters of given size and its standard deviation
where (presenting the following transformations may be useful for clarity as it was not presented explicitly before)

### v2

@@ A.2 Generating function method
Below, we will show how to obtain Eq. (12) from Eq. (13). This derivation Lepek_2019 is worth reviewing because of (at least) two important reasons. First, it proves that Eqs. (12) and (13) define the same sequence. The recursive form will be later used to obtain $x_{g}$ for any given kernel. Second, the methodology to obtain Eq. (12) from Eq. (13), i.e., the generating function method, was the main mathematical tool applied to derive explicit solutions to several important kernels (e.g., product, additive, condensation, electrorheological).
Expanding binomials in Eq. (13) we have
@@
Transforming left-hand side gives
@@
$G(z)\equiv\sum^{\infty}_{g=1}{y_{g}z^{g}}$ (120)
is the generating function for $y_{g}$. Analogously, after transforming right-hand side of Eq. (118) we obtain
@@
By the above transformations, Eqs. (119) and (121), from Eq. (118) we obtained differential equation for $G(z)$,
$z\frac{\partial}{\partial z}G(z)-G(z)=\frac{1}{2}{\left(G(z)\right)}^{2}.$ (122)
As Eq. (122) is an ODE which can be solved by separating variables, namely,
$\frac{{d}G(z)}{\frac{1}{2}G^{2}(z)+G(z)}=\frac{{d}z}{z},$ (123)
its solution may be written as
$\begin{split}G(z)&=\frac{2Cz}{1-Az}=2Cz\left(\frac{1}{1-Cz}\right)\\ &=2Cz\sum^{\infty}_{g=0}{{\left(Cz\right)}^{g}}=2\sum^{\infty}_{g=1}{{\left(Cz\right)}^{g}}\end{split}$ (124)
with $C$ being a constant.
@@
$2C^{g}=\frac{x_{g}}{g!\left(g-1\right)!}.$ (125)
Assuming reasonably that $x_{1}=1$, we can calculate the constant $C=1/2$ to obtain
@@
A.3 Transformations for standard deviation
One can calculate standard deviation corresponding to the average number of clusters of a given size as Fronczak_2018
@@ A.3 Transformations for standard deviation
where (presenting the following transformations may be useful for clarity as they were not presented explicitly before)